OpenStudy (tinash):

Patterns etc. Please help. Need to get this done and am not sure if it is right.

OpenStudy (tinash):

Q5 (C) am not sure about it

OpenStudy (anonymous):

What grade is this?

OpenStudy (phi):

yes, it looks good

OpenStudy (tinash):

11th grade. How do I do (c) though?

OpenStudy (phi):

look at the first two rows in your table hour 1 2 3 4 5 white 2 4 6 8 10 you need to find a pattern.

OpenStudy (tinash):

I know but the formula ... I don't know if I got it or not. Tn=a+(n-1)d Tn=2+(n-1)2 2+5n-2 5n I don't know if I am right or wrong

OpenStudy (phi):

The first thing people try is find the difference between the values the difference in the white is 2 and the difference in the hours is 1 if you think of them as (x,y) points, you will see that you have a *line* with slope 2. If you know about point-slope (for a line) you could write w-2 = 2(h-1) which simplifies to w-2 = 2h-2 w= 2h

OpenStudy (tinash):

No I don't know anything about that...

OpenStudy (phi):

ok, you can also use your formula you are ok up to Tn=2+(n-1)2 now "distribute" the 2 what do you get ?

OpenStudy (tinash):

I have no clue....

OpenStudy (tinash):

I am really sorry am trying I just don't now how to distribute it

OpenStudy (phi):

focus on the (n-1)2 that means 2 times (n-1) think of the parens as a package with n and -1 in it, and you have two packages. so how many n's do you have and how many -1's ?

OpenStudy (tinash):

2n-1

OpenStudy (phi):

ok on the 2n but if you have 2 packages, and each one has a -1, don't you have 2 of the -1's also ?

OpenStudy (tinash):

oh wait 2n-2

OpenStudy (phi):

yes that is "distributing"

OpenStudy (phi):

Tn=2+(n-1)2 becomes Tn= 2 +2n -2

OpenStudy (phi):

you can simplify the +2 -2 to zero

OpenStudy (tinash):

so it is just 2n?

OpenStudy (phi):

yes, the rule is Wh = 2h (I used W for "white tiles" and h for hour) if we check your table, at h=3 (for example), the formula says W= 2*3= 6 we should have 6 white tiles. and it does. The formula works.

OpenStudy (phi):

btw, I should write W=2h not Wh (because Wh might mean W*h, which is not what we want) I meant \[ W_h= 2h \] which is a fancy way to say, the number of white tiles for "h" hours is 2*h

OpenStudy (tinash):

I got it .. thank you <3

OpenStudy (phi):

but I think they want you to use \(T_n\), so I would use that \[ T_n= 2n\]

OpenStudy (tinash):

Yep c: merci

OpenStudy (phi):

The black tiles are even simpler Tn= 2+ n

OpenStudy (phi):

For the total, add the two formulas can you do that ?

OpenStudy (tinash):

Nope never done it

OpenStudy (phi):

write down the number of white tiles (i.e. write 2n) then put a + sign then write down the number of black tiles what does that look like ?

OpenStudy (tinash):

simultaneous equation? or I think linear

OpenStudy (phi):

No, I meant just write down the number of white tiles (i.e. write 2n) then put a + sign then write down the number of black tiles

OpenStudy (tinash):

2n + tn=2+n

OpenStudy (phi):

The formula for the white tiles is 2n (think of that as a number) and the number of black tiles is 2+n to show those being added, you write them next to each other, and put + in between them can you try to do that ?

OpenStudy (tinash):

yes I have done that.

OpenStudy (phi):

can you type it here?

OpenStudy (tinash):

2n+2+n

OpenStudy (phi):

yes, and that is the answer. But most people would simplify it. you have 2 n's plus another n. how many n's do you have

OpenStudy (tinash):

3n

OpenStudy (phi):

yes, and your formula 2n+2+n simplifies to 3n + 2

OpenStudy (tinash):

THANK YOU